a metal rod with density od 7000kg/m3 and diameter 40mm has a mass if 1.76kg. calculate its length
1
Expert's answer
2012-10-03T12:05:05-0400
Let's make following denotations:

Inputs are inconsistent, task is not clear

a metal rod with density od 7000kg/m³ and diameter 40mm has a mass if 1.76kg. calculate its length

The volume of a rod is

V = M/p = 1.76[kg]/7000[kg/m³].

From the other side,

V = πR²L ==> L = V/(πR²) = V/(π(D/2)²) = (M/p)/(π(D/2)²) = 4M/(pπD²) = 4*1.76[kg]/(7000kg/m³*3.14*0.04[m]) ≈ 0.008 m.

So, the length of the rod is approximately 0.8 cm.